OpenStudy (anonymous):

Help me with this i'm trying to learn I'm grade 5 5 wholes 5 out of 6 divided by 7

OpenStudy (anonymous):

can any one help

zepdrix (zepdrix):

\[\huge 5\frac{ 5 }{ 6 }\div 7\] Is this what the problem looks like?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\large 5\frac{ 5 }{ 6 }=5\frac{ 6 }{ 6 }+\frac{ 5 }{ 6 }\]\[\large =\frac{ 30 }{ 5 }+\frac{ 5 }{ 6 }\]\[\large =\frac{ 35 }{ 6 }\] Do you understand how to change a mixed number into an improper fraction like this? :)

zepdrix (zepdrix):

Oooo cool fire bird! :Ooo

OpenStudy (anonymous):

thanks